OpenStudy (anonymous):

A child drops his toy car from a 2nd floor window that is 7.3 meters high. Find the velocity of the toy car just before it hits the ground below. Ignore any affects of air resistance. **not sure how to do this one! :/ thank you:)

OpenStudy (phi):

I would write down the formulas you know... (I do that to waste time until a good thought happens...)

OpenStudy (anonymous):

haha okie! so f=m*a v(t)=v0+a*t v(t)=a*t ermm not sure what other ones would apply here?

OpenStudy (phi):

the distance = 1/2 a t^2

OpenStudy (anonymous):

ohh so is that what i use for this problem?

OpenStudy (phi):

also a= g = 9.8 m/s or 32 ft/sec

OpenStudy (phi):

next, write down the numbers (variables that have know values) which equations (if any) can be solved for an unknown variable? if we do solve, will that be helpful ?

OpenStudy (anonymous):

so 7.3=height? and final velocity would be 0 ?

OpenStudy (anonymous):

@phi?

OpenStudy (michele_laino):

do you know the Torricelli formula? @iheartfood

OpenStudy (anonymous):

yes, but i forget which equation that is:/

OpenStudy (phi):

7.3 m = 1/2 9.8 m/s^2 t^2 solve for t

OpenStudy (phi):

after you find t, use v= 9.8 t to find the velocity.

OpenStudy (anonymous):

oh, so 7.3=4.9 * t^2 2.4=t^2 ? t=1.55? @phi?

OpenStudy (phi):

7.3=4.9 * t^2 t^2= 7.3/4.9 = ?

OpenStudy (anonymous):

t^2= 1.489 = 1.49 ?

OpenStudy (anonymous):

t=1.22 ? @phi?

OpenStudy (phi):

yes, that looks better. now you can find v , right ?

OpenStudy (anonymous):

is it v=a*t ?

OpenStudy (anonymous):

v= 9.8 and t=1.22 ? solve for a? or is something else? :/

OpenStudy (phi):

a is acceleration (here due to gravity) and we know a. (it's "g" =9.8 m/s^2 ) we don't (yet) know v after falling for 1.22 seconds

OpenStudy (anonymous):

ohh okay! so v=9.8 * 1.22 v=11.956 m/s ?

OpenStudy (phi):

but put in the units

OpenStudy (anonymous):

oh like v= 9.8 m/s^2 * 1.22 s = 11.956 m/s ?

OpenStudy (phi):

\[ v=9.8 \frac{m}{s^\cancel2} 1.22 \cancel{s} = 11.956 \frac{m}{s}\] after we "cancel" the seconds, the final units work out to m/s. This shows we are probably correct.. we got the expected units.

OpenStudy (anonymous):

ah yes!! :) yay thank you so much @phi :)

OpenStudy (phi):

yw